Treasury bills are currently paying 5 percent and the inflation rate is 2.7 percent.

a.

What is the approximate real rate of interest? (Enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)

b. What is the exact real rate? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
Answer #1

a.approximate real rate of interest=nominal rate-inflation rate

=5-2.7

=2.3%

b. exact real rate=(1+nominal rate)/(1+inflation rate)-1

=[(1+0.05)/(1+0.027)]-1

=(1.05/1.027)-1

=2.24%(Approx).
Answer #2

ANSWER 


a.


Approximate real rate

= Treasury rate - Inflation rate

= 5% - 2.7%

= 2.3% (ANSWER).


b.


Exact real rate 

= (1 +Treasury rate) / (1 + inflation rate) - 1

= (1 + 0.05) / (1 + 0.027) - 1

= 0.0224

= 2.24%  (ANSWER).


answered by: Tulsiram Garg